Stephen purchased a diamond engagement ring for $3,000. His down payment was $1700, and he made 24 monthly payments of $59.99. Find the APR
Answer by stanbon(75887)

Stephen purchased a diamond engagement ring for $3,000. His down payment was $1700, and he made 24 monthly payments of $59.99. Find the APR
-------
3000-1700 1300
24*59.99 = 1439.76
-----
Interest paid over 2 years = 439.76
------
I = P*r*t
439.76 = 1439.76*r*2
---------
2r = 0.3054
Rate = 0.1527 = 15.27%
---------------
Cheers,
Stan H.
-----------